40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 510     Entry time: Thu Feb 24 01:40:00 2011
Author: Frank 
Type: DailyProgress 
Category: BEAT 
Subject: influence of power fluctuations 

I had a closer look into effects caused by power fluctuations. To summarize : It is easy to measure TFs from power fluctuations to any other point in the system. SNR is good to very good.

I mainly focused on the effect of changes in shape of the TF from power fluctuations into changes of the beat frequency when changing the power levels anywhere in the system, especially in the cavity paths.
After playing a couple of hours and getting a feeling for what's going on i finally realized that only if changing the power in the ACAV-path strange things happen. This includes changing the total power going to both cavities as well. As taking TF's is taking too long at low frequencies i decided to switch from swept sine measurements to a simple digital modulation of the laser power with 100ms period. This is slow enough to see thermal effects in the cavity and other strange things which i will show below.

For the first set of measurements i reduced the power to the REFCAV path to about 460uW. The total modulation is very large, about 15.6%. The max power to ACAV was about 5.76mW, which i reduced later.
The following graphs shows the response of the beat signal to the digital modulation, beat signal calibrated to Hz, modulation signal a.u.., but the amplitude didn't change between measurements.

 powermod_ACAV.jpg

 on the left graph (measured at very low power) one can see that the power modulation causes a sudden changes in frequency of the beat signal.
This can be caused by everything producing an offset in the EP which depends on power, e.g. RF-AM, higher-order modes in reflection etc. Nothing unusual so far.
 If the optical power going to the ACAV is increased one can find a point where the change in frequency is almost gone, showed in the center graph.
When the power in increased further the sign of the frequency shift changes! Now some thermal effect becomes dominant, most likely the cavity as the period is 10s and after 5s equilibrium isn't reached.

Now this explains everything i saw before: When measuring the TF from power fluctuations to frequency shift and varying the power the shape of the TF changed because we have two different effects which can even cancel at the right power level send to the cavity!

 

As we know already that the size of the error signals is too tiny this effect will probably be almost gone once we fixed the demodulation for ACAV. I tried to reduce it a little bit by re-aligning the cavity but no luck - it's already aligned very well. I didn't try the EOM asthis changes the alignment to the other cavity as well and i was to lazy to re-align everything

ELOG V3.1.3-